the initial velocity of the partical is u (at t=0) and acceleration f=at( here t = time and a is constant).which ralation is right=
1.v=u+at
.v=u+at*t/2

Grade:12

1 Answers

Dhruvit Raithatha
6 years ago
f = a\cdot t \\ \int f = v = \int at\cdot dt\\ v = \frac{at^2}{2} + C\\ \text{Therefore, final velocity} = u +\frac{at^2}{2}